# Local structural disorder in crystalline materials

### v1

@@ 1 Introduction
Local disorder is an intrinsic feature of soft, anharmonic materials, first realized in the late 60’s by Comes et al. for oxide perovskites 1. While standard diffraction analyses of crystalline materials suggest high-symmetry monomorphous structures, local fluctuations in atomic positions give rise to a rich configurational landscape, as suggested by more advanced experimental approaches 2, 3, 4, 5, 6, 7, 8, 9. These variations, linked to configurational entropy, are termed positional polymorphism 10 and impact key physical properties in ways that are not captured by conventional, monomorphous models. Some communities refer to this effect as static disorder or split-site occupancy, terms that are related but not strictly identical. Recent advances in first-principles theory have enabled the explicit treatment of such disorder 10, 11, offering a new lens to interpret and predict material behavior in energy applications.
@@
Local disorder is a typical characteristic of room temperature phases of materials, and particular of anharmonic materials featuring a multiwell potential energy surface (PES). The high symmetry configuration corresponds to a local maximum on the PES, and locally disordered configurations correspond to energetically favorable minima, as shown in Figure 1a. In both cases, the system corresponds to a static equilibrium state at $T=0$ K. At higher temperatures, ultraslow relaxation occurs between locally disordered configurations. The polymorphous framework does not explicitly model these dynamics, but instead represents structural disorder through optimized supercells or a limited sampling of them. This raises a fundamental question: which structural configurations should be considered as physically realistic for ab initio electronic structure calculations, particularly those based on density functional theory (DFT)? This choice also has profound implications for the coupling of carriers with lattice dynamics, i.e. electron-phonon coupling 12, as well as for phonon-phonon scattering which determine key optical and transport properties that govern the performance of energy devices.
@@
In this Perspective, we present compelling evidence for the necessity of using locally disordered structures to achieve accurate predictions of electronic structure, vibrational dynamics, and electron-phonon interactions. Such an approach is crucial for capturing temperature-dependent properties, including vibrational spectra, effective masses, band gaps, and carrier mobilities, where theory results demonstrate excellent agreement with experimental data. We provide an analysis of ongoing theoretical and experimental efforts addressing the role of local disorder in materials, as well as outline limitations of the framework and emerging future directions. We discuss how local disorder can lead to improved predictions of key phenomena such carrier and thermal transport as well as polaron formation. We believe this Perspective will contribute to a deeper understanding of the intricate interplay between structural disorder, electronic structure, anharmonicity, and electron-phonon coupling, while also accelerating the development of more accurate theoretical frameworks for predicting energy-relevant properties of functional materials. Among the numerous materials known to feature local disorder 1, 13, 14, 15, 16, 17, 9, 18, emphasis is given to halide perovskites 2, 10, 19, a rapidly advancing area in energy research.
@@
Figure 1: Impact of local disorder on the electronic structure. (a) Schematic representation of the PES illustrating the high symmetry and locally disordered structures, and the corresponding energy lowering $\Delta U$. (b) Locally disordered (polymorphous) structures of cubic CsPbI3 and CsSnI3 obtained by the ASDM using a $4\times 4\times 4$ supercell and DFT geometry optimization. (c) Average DFT charge densities of high symmetry and locally disordered cubic CsSnI3. (d) DFT electron spectral function of polymorphous CsSnI3 calculated by band structure unfolding 20. Data are from Refs. 11, 21
@@ 2 Electronic structure
Local disorder in materials refers to spatially correlated deviations of atoms from their high symmetry positions, which on average preserve the crystallographic space group symmetry of the structure 10. To simulate positional polymorphism, a common approach is to start from a high symmetry supercell structure and introduce atomic displacements via random nudges 10 or special displacements along phonons 20, 11. The structure is then relaxed with the lattice constants fixed. This method enables the system to explore symmetry-breaking domains and identify energetically favorable minima in the PES, as illustrated in Figure 1a. The energy lowering with respect to the monomorphous structure is indicated as $\Delta U$.
@@
Focusing on non-polar fluctuations, the degree of local disorder is quantified by calculating the deviation of the B-X-B angle ($\theta_{\rm B-X-B}$) from the ideal value of 180 in the cubic structure 21. In the examples shown in Figure 1b, local disorder is greater in CsPbI3 ($\Delta\theta_{\rm B-X-B}$ = 26.7) than in CsSnI3 ($\Delta\theta_{\rm B-X-B}$ = 23.2), which is also reflected in the corresponding $\Delta U$ values for each structure. Due to stronger orbital overlap and enhanced bond covalency, CsPbI3 has a larger lattice constant (6.25 Å vs. 6.14 Å in CsSnI3), which in turn provides greater structural flexibility and a higher tendency toward local symmetry breaking. As discussed in Ref. 19, another key parameter determining the degree of local disorder in cubic halide perovskites is the tolerance factor. This is strongly influenced by the size of the A-site cation, with higher values indicating less local distortions. For example, FA-based compounds have been shown to exhibit the lowest degree of positional polymorphism due to larger tolerance factors than MA-based and Cs-based compounds 21. Aziridinium (AZ)-based halide perovskites, a recent addition to the family of A-site cations 23, 24, are expected to exhibit a degree of local disorder in the range ($\Delta\theta_{\rm B-X-B}$ = 13–18) higher than FA-based compounds ($\Delta\theta_{\rm B-X-B}$ = 9–12), yet slightly lower than that typically observed in MA-based systems ($\Delta\theta_{\rm B-X-B}$ = 14–1919. This is because the estimated ionic radius of AZ is approximately 2.30 Å, which is intermediate between FA (2.53 Å) and MA (2.17 Å); the degree of local disorder in AZ-based compounds remains to be confirmed by first-principles calculations.
@@
Figure 1c shows that using a monomorphous model in DFT calculations for cubic CsSnI3 and CsPbI3, including the effect of spin-orbit coupling (SOC), leads to an unphysical result. The charge density at the valence band maximum (VBM) and conduction band minimum (CBM) exhibits a spurious exchange in orbital character, with the VBM formed primarily by B-site metal p orbitals and the CBM showing contributions from X-site halogen p states and B-site s states. This results to a metalic-like behavior and negative electron effective masses at the R-point for CsSnI3 as shown in Figure 1d. In contrast to common beliefs, this is not only a deficiency of DFT but a consequence of using the incorrect structural model. As shown in Figure 1c for the average polymorphous structure, the expected orbital character in the charge density at the band extrema is recovered, leading to an overall band gap opening of 0.56 eV (Figure 1d). Therefore, the apparent failure to describe the semiconducting behavior often arise from neglecting local symmetry breaking rather than from the intrinsic inability of DFT to handle electron correlations. This has been also observed, e.g., in oxide perovskites 14, low-dimensional quantum materials 17, transition metal monoxides 13, transition metal chalcogenides 25, and Cu-based superionic conductors 16. We stress that while DFT combined with realistic structures captures key qualitative features, accurate band gaps and effective masses require advanced methods like hybrid functionals or $GW$ to account more accurately for electron correlation 26, 21.
@@ 3 Structural and Vibrational properties
Figure 2: Impact of local disorder on structural and vibrational properties. (a) Calculated XRD patterns of monomorphous and polymorphous cubic SrTiO3 compared to experiment. Reproduced from Ref. 15, copyright (2021), with permission from Elsevier. (b) Calculated PDF of monomorphous and polymorphous cubic MAPbI3 compared with experiment. Reproduced with permission from Ref. 10, copyright (2020) by the American Physical Society. (c) Phonon spectral function of polymorphous CsPbI3 obtained by phonon unfolding. Data extracted from Ref. 11. (d) Phonon energies in reciprocal space measured for MAPbBr3 with inelastic neutron scattering. Figure adapted from Ref. 27 under $C$ BY 4.0. (e) Calculated Raman spectra of CsPbBr3 by MD compared to experimental measurements. Reproduced with permission from Ref. 28, copyright (2017) by the American Physical Society. (f) Calculated phonon DOS of monomorphous and polymorphous cubic Cu2Se compared with experiment. Data from Ref. 16.
@@
Positional polymorphism also strongly affects the phonon spectra, leading to strongly overdamped vibrational dynamics, as shown for CsPbI3 by Zacharias et al. 11 (color map in Figure 2c). Employing a monomorphous structure together with the harmonic approximation to compute phonons results in pronounced dynamical instabilities in the phonon dispersion, as shown by the black lines in Figure 2c. This occurs because the monomorphous configuration corresponds to a local maximum on the PES, leading to negative curvatures along certain vibrational modes, manifested as imaginary phonon frequencies. Instead, using a polymorphous network yields dynamically stable phonons, with acoustic branches emerging near the $\Gamma$-point from a background of nearly dispersionless optical vibrations, consistent with inelastic neutron scattering measurements in halide perovskites 27, shown in Figure 2d for MAPbBr3. The strongly overdamped vibrations indicate a breakdown of the conventional phonon quasiparticle picture that is typically assumed in the monomorphous model. This breakdown becomes even more pronounced in hybrid halide perovskites due to enhanced mode mixing between polar optical phonons and the internal vibrations of the MA or FA molecules 21. Raman scattering measurements on halide perovskites reported by Yaffe et al. 28 reveal strongly correlated vibrational dynamics, as evidenced by a pronounced central peak spanning a broad frequency range, as shown in Figure 2e. This feature is absent in monomorphous harmonic systems, which typically exhibit sharp and well-defined Raman peaks. Interestingly, ab initio molecular dynamics (MD) can reproduce the central peak of the high temperature phase of CsPbBr3, demonstrating that MD can capture both thermal vibrations and local disorder effects. Similarly to Raman spectra, another observable that can be used to identify strongly overdamped, anharmonic vibrations is the phonon density of states (DOS). Figure 2e shows the phonon DOS of Cu2Se, a polymorphous material with high degree of local disorder, as calculated by Wang et al. 16 using the anharmonic special displacement method (ASDM) 20, 29. The presence of positional polymorphism leads to significant spectral broadening, in good agreement with experiment 30, and contrasts with the well-defined peaks calculated for the monomorphous structure.
@@
The breakdown in the phonon quasiparticle picture, apart from resulting in broadened phonon spectra and reduced lifetimes, it also leads to enhanced phonon-phonon scattering, suppressing lattice thermal conductivity due to increased energy dissipation and reduced phonon mean free paths 16, 9. For thermoelectric materials, this suppression of thermal conductivity is beneficial, as it enhances the figure of merit, provided that the electronic transport remains relatively unaffected. Therefore, understanding and controlling disorder-induced phonon scattering is not only crucial for accurately modeling thermal transport but also offers a strategy to engineer high-performance thermoelectrics.
@@ 4 Thermal fluctuations in an anharmonic lattice
In Figures 3a and b, we illustrate the two sources of anharmonicity that influence the vibrational dynamics: one is positional polymorphism, as discussed for Figure 2, and the other arises from temperature-dependent phonon self-energy corrections. In the context of lattice dynamics, the loop diagram originating from quartic anharmonicity (Figure 3b) represents the first correction to phonon frequencies beyond the harmonic approximation 31. It corresponds to a process where a phonon interacts with the thermal fluctuations of the lattice, mediated by a virtual phonon through a quartic vertex. Although quartic terms are associated with the 4th-order potential, the loop diagram does not represent a real scattering process among four phonons. Instead, it captures a self-interaction mechanism, where thermal fluctuations in the lattice feed back into the mode’s own dynamics. This correction affects only the real part of the phonon self-energy and introduces no temperature-dependent broadening or lifetime effects, usually obtained by the so called bubble diagram 31. Within the self-consistent phonon (SCP) theory, temperature-dependent phonons are computed by iteratively solving an effective Hessian derived from interatomic forces. To obtain the final renormalized phonon frequencies and eigenvectors, SCP theory effectively resums anharmonic loop contributions to infinite order, offering a physically grounded and non-perturbative description of strong anharmonicity. There are several state-of-the-art methods and codes that implement the SCP theory and extensions beyond it, including the Temperature Dependent Effective Potential (TDEP32, ALAMODE 31, stochastic self-consistent harmonic approximation (SSCHA33, and the ASDM 29. Within the same level of approximation, such as consistent calculation settings and order of anharmonicity, these methods are expected to yield identical results.
@@
Figure 3: Impact of local disorder on electron-phonon coupling. (a,b,c) Locally disordered structure and diagrammatic representation of loop diagram contributing to temperature-dependent anharmonicity and the basic electron-phonon coupling process. Lines represent electrons with wavector $\bf k$} and $\bf k+q$} and the zigzag lines the anharmonic phonon. (d,e) Calculated band gap renormalization of monomorphous and polymorphous cubic CsPbI3 and tetragonal/cubic CsPbBr3 by ASDM compared with experiment. Data taken from Refs. 21, 11. (f) Calculated electron and hole effective mass of monomorphous and polymorphous cubic CsPbCl3. Data extracted from Ref. 21. Vertical dashed lines indicate temperatures where phase transitions take place.

### v2

# Local Structural Sisorder in Crystalline Materials

@@ Abstract
Local disorder is an intrinsic feature of soft, anharmonic materials, first realized in the late 60’s by Comes et al. for oxide perovskites 1. While standard diffraction analyses of crystalline materials suggest high-symmetry monomorphous structures, local fluctuations in atomic positions give rise to a rich configurational landscape, as suggested by more advanced experimental approaches 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12. These variations, linked to configurational entropy, are termed positional polymorphism 13 and impact key physical properties in ways that are not captured by conventional, monomorphous models. Some communities refer to this effect as static disorder or split-site occupancy, terms that are related but not strictly identical. Recent advances in first-principles theory have enabled the explicit treatment of such disorder 13, 14, offering a new lens to interpret and predict material behavior in energy applications.
@@
Local disorder is a typical characteristic of room temperature phases of materials, and particular of anharmonic materials featuring a multiwell potential energy surface (PES). The high symmetry configuration corresponds to a local maximum on the PES, and locally disordered configurations correspond to energetically favorable minima, as shown in Figure 1a. At $T=0$ K, both cases correspond to static equilibrium states. At higher temperatures, ultraslow relaxation occurs between locally disordered configurations. The polymorphous framework does not explicitly model these dynamics, but instead represents structural disorder through optimized supercells or a limited sampling of them. This raises a fundamental question: which structural configurations should be considered as physically realistic for ab initio electronic structure calculations, particularly those based on density functional theory (DFT)? This choice also has profound implications for the coupling of carriers with lattice dynamics, i.e. electron-phonon coupling 15, as well as for phonon-phonon scattering which determine key optical and transport properties that govern the performance of energy devices.
@@
In this Perspective, we present compelling evidence for the necessity of using locally disordered structures to achieve accurate predictions of electronic structure, vibrational dynamics, and electron-phonon interactions. Such an approach is crucial for capturing temperature-dependent properties, including vibrational spectra, effective masses, band gaps, and carrier mobilities, where theory results demonstrate excellent agreement with experimental data. We provide an analysis of ongoing theoretical and experimental efforts addressing the role of local disorder in materials, as well as outline limitations of the framework and emerging future directions. We discuss how local disorder can lead to improved predictions of key phenomena such carrier and thermal transport as well as polaron formation. We believe this Perspective will contribute to a deeper understanding of the intricate interplay between structural disorder, electronic structure, anharmonicity, and electron-phonon coupling, while also accelerating the development of more accurate theoretical frameworks for predicting energy-relevant properties of functional materials. Among the numerous materials known to feature local disorder 1, 16, 17, 18, 19, 20, 10, 21, emphasis is given to halide perovskites 2, 13, 22, a rapidly advancing area in energy research.
@@
Figure 1: Impact of local disorder on the electronic structure. (a) Schematic representation of the PES illustrating the high symmetry and locally disordered structures, and the corresponding energy lowering $\Delta U$. (b) Locally disordered (polymorphous) structures of cubic CsPbI3 and CsSnI3 obtained by the ASDM using a $4\times 4\times 4$ supercell and DFT geometry optimization. (c) Average DFT charge densities of high symmetry and locally disordered cubic CsSnI3. (d) DFT electron spectral function of polymorphous CsSnI3 calculated by band structure unfolding 23. The band structure of the monomorphous model is shown as red. The band gap opening due to local disorder is $\Delta E_{\rm g}=0.56$ eV. Data are from Refs. 14, 24
@@ 1 Electronic structure
Local disorder in materials refers to spatially correlated deviations of atoms from their high symmetry positions, which on average preserve the crystallographic space group symmetry of the structure 13. To simulate positional polymorphism, a natural approach is to employ ab initio molecular dynamics (MD) with extended equilibration and production times to sample the relevant configurational space. Another approach is to start from a high symmetry supercell structure and introduce atomic displacements via random nudges 13 or special displacements along phonons 23, 14. The structure is then relaxed with the lattice constants fixed. This method enables the system to explore symmetry-breaking domains and identify energetically favorable minima in the PES, as illustrated in Figure 1a. The energy lowering with respect to the monomorphous structure is indicated as $\Delta U$.
@@
Focusing on non-polar fluctuations, the degree of local disorder is quantified by calculating the deviation of the B-X-B angle ($\theta_{\rm B-X-B}$) from the ideal value of 180 in the cubic structure 24. In the examples shown in Figure 1b, local disorder is greater in CsPbI3 ($\Delta\theta_{\rm B-X-B}$ = 26.7) than in CsSnI3 ($\Delta\theta_{\rm B-X-B}$ = 23.2), which is also reflected in the corresponding $\Delta U$ values for each structure. Due to stronger orbital overlap and enhanced bond covalency, CsPbI3 has a larger lattice constant (6.25 Å vs. 6.14 Å in CsSnI3), which in turn provides greater structural flexibility and a higher tendency toward local symmetry breaking. As discussed in Ref. 22, another key parameter determining the degree of local disorder in cubic halide perovskites is the tolerance factor. This is strongly influenced by the size of the A-site cation, with higher values indicating less local distortions. For example, FA-based compounds have been shown to exhibit the lowest degree of positional polymorphism due to larger tolerance factors than MA-based and Cs-based compounds 24. Aziridinium (AZ)-based halide perovskites, a recent addition to the family of A-site cations 26, 27, are expected to exhibit a degree of local disorder in the range ($\Delta\theta_{\rm B-X-B}$ = 13–18) higher than FA-based compounds ($\Delta\theta_{\rm B-X-B}$ = 9–12), yet slightly lower than that typically observed in MA-based systems ($\Delta\theta_{\rm B-X-B}$ = 14–1922. This is because the estimated ionic radius of AZ is approximately 2.30 Å, which is intermediate between FA (2.53 Å) and MA (2.17 Å); the degree of local disorder in AZ-based compounds remains to be confirmed by first-principles calculations.
@@
Figure 1c shows that using a monomorphous model in DFT calculations for cubic CsSnI3 and CsPbI3, including the effect of spin-orbit coupling (SOC), leads to an unphysical result. The charge density at the valence band maximum (VBM) and conduction band minimum (CBM) exhibits a spurious exchange in orbital character, with the VBM formed primarily by B-site metal p orbitals and the CBM showing contributions from X-site halogen p states and B-site s states. This results in a spurious negative gap of -0.27 eV and negative electron effective masses at the R-point for CsSnI3 as shown in Figure 1d. In contrast to common beliefs, this is not only a deficiency of DFT to handle electron correlations but a consequence of using the incorrect structural model without local disorder. As shown in Figure 1c for the average polymorphous structure, the expected orbital character in the charge density at the band extrema is recovered, leading to an overall band gap opening of 0.56 eV (Figure 1d). This value correlates with the B-X-B bond angle changes associated with local disorder and tolerance factor, consistent with trends calculated for other halide perovskites 22. Accounting for local symmetry breaking is essential to recover reliable band gap behavior, as also shown for oxide perovskites 17, low-dimensional quantum materials 20, transition metal monoxides 16, transition metal chalcogenides 28, and Cu-based superionic conductors 19. We stress that while DFT combined with locally disordered structures captures key qualitative features, accurate band gaps and effective masses require advanced methods like hybrid functionals or $GW$ to account more accurately for electron correlation 29, 24.
@@ 2 Structural and Vibrational properties
Figure 2: Impact of local disorder on structural and vibrational properties. (a) Calculated XRD patterns of monomorphous and polymorphous cubic SrTiO3 compared to experiment. Reproduced from Ref. 18, copyright (2021), with permission from Elsevier. (b) Calculated PDF of monomorphous and polymorphous cubic MAPbI3 compared with experiment. Reproduced with permission from Ref. 13, copyright (2020) by the American Physical Society. (c) Phonon spectral function of polymorphous CsPbI3 obtained by phonon unfolding. Data extracted from Ref. 14. (d) Phonon energies in reciprocal space measured for MAPbBr3 with inelastic neutron scattering. Figure adapted from Ref. 30 under $C$ BY 4.0. (e) Calculated Raman spectra of CsPbBr3 by MD compared to experimental measurements. Reproduced with permission from Ref. 31, copyright (2017) by the American Physical Society. (f) Calculated phonon DOS of monomorphous and polymorphous cubic Cu2Se compared with experiment. Data from Ref. 19.
@@
Positional polymorphism also strongly affects the phonon spectra, leading to strongly overdamped vibrational dynamics, as shown for CsPbI3 by Zacharias et al. 14 (color map in Figure 2c). Employing a monomorphous structure together with the harmonic approximation to compute phonons results in pronounced dynamical instabilities in the phonon dispersion, as shown by the black lines in Figure 2c. This occurs because the monomorphous configuration corresponds to a local maximum on the PES, leading to negative curvatures along certain vibrational modes, manifested as imaginary phonon frequencies. Instead, using a polymorphous network yields dynamically stable phonons, with acoustic branches emerging near the $\Gamma$-point from a background of nearly dispersionless optical vibrations, consistent with inelastic neutron scattering measurements in halide perovskites 30, shown in Figure 2d for MAPbBr3. The strongly overdamped vibrations lead to the breakdown of the conventional phonon quasiparticle picture that is typically assumed in the monomorphous model. This breakdown becomes even more pronounced in hybrid halide perovskites due to enhanced mode mixing between polar optical phonons and the internal vibrations of the MA or FA molecules 24. We remark that in many systems, unstable modes become anharmonically stabilised at higher temperatures, still retaining a well-defined quasiparticle picture as shown by MD-based TDEP calculations 32.
Raman scattering measurements on halide perovskites reported by Yaffe et al. 31 reveal strongly correlated vibrational dynamics, as evidenced by a pronounced central peak spanning a broad frequency range, as shown in Figure 2e. This feature is absent in monomorphous harmonic systems, which typically exhibit sharp and well-defined Raman peaks. Interestingly, MD can reproduce the central peak of the high temperature phase of CsPbBr3, demonstrating that MD can capture both thermal vibrations and local disorder effects. Furthermore, MD has been proven powerful to capture overdamped phonon spectra in halide perovskites by analysis of velocity autocorrelation functions 33. Another observable that can be used to identify strongly overdamped, anharmonic vibrations is the phonon density of states (DOS). Figure 2e shows the phonon DOS of Cu2Se, a polymorphous material with high degree of local disorder, as calculated by Wang et al. 19 using the anharmonic special displacement method (ASDM) 23, 34. The presence of positional polymorphism leads to significant spectral broadening, in good agreement with experiment 35, and contrasts with the well-defined peaks calculated for the monomorphous structure. This behavior is consistent with prior observations of strongly overdamped vibrational spectra in superionic materials 36. Oxide perovskites represent another class of materials where overdamped phonon dynamics have been measured and supported through MD simulations 37.
@@
The breakdown in the phonon quasiparticle picture, apart from resulting in broadened phonon spectra and reduced lifetimes, it also leads to enhanced phonon-phonon scattering, suppressing lattice thermal conductivity due to increased energy dissipation and reduced phonon mean free paths 38, 36, 19, 10. For thermoelectric materials, this suppression of thermal conductivity is beneficial, as it enhances the figure of merit, provided that the electronic transport remains relatively unaffected. Therefore, understanding and controlling disorder-induced phonon scattering is not only crucial for accurately modeling thermal transport but also offers a strategy to engineer high-performance thermoelectrics.
@@ 3 Thermal fluctuations in an anharmonic lattice
In Figures 3a and b, we illustrate the two sources of anharmonicity that influence the vibrational dynamics: one is positional polymorphism, as discussed for Figure 2, and the other arises from temperature-dependent phonon self-energy corrections. In the context of lattice dynamics, the loop diagram originating from quartic anharmonicity (Figure 3b) represents the first correction to phonon frequencies beyond the harmonic approximation 39. It corresponds to a process where a phonon interacts with the thermal fluctuations of the lattice, mediated by a virtual phonon through a quartic vertex. Although quartic terms are associated with the 4th-order potential, the loop diagram does not represent a real scattering process among four phonons. Instead, it captures a self-interaction mechanism, where thermal fluctuations in the lattice feed back into the mode’s own dynamics. This correction affects only the real part of the phonon self-energy and introduces no temperature-dependent broadening or lifetime effects, usually obtained by the so called bubble diagram 39. Within the self-consistent phonon (SCP) theory, temperature-dependent phonons are computed by iteratively solving an effective Hessian derived from interatomic forces. To obtain the final renormalized phonon frequencies and eigenvectors, SCP theory effectively resums anharmonic loop contributions to infinite order, offering a physically grounded and non-perturbative description of strong anharmonicity.
There are several state-of-the-art methods and codes that implement the SCP theory and extensions beyond it, including the Temperature Dependent Effective Potential (TDEP40, ALAMODE 39, stochastic self-consistent harmonic approximation (SSCHA41, and the ASDM 34. Within the same level of approximation, such as consistent calculation settings and order of anharmonicity, these methods are expected to yield identical results. We emphasize that these approaches to temperature-dependent anharmonicity are conceptually different than the polymorphous framework. They primarily incorporate phonon self-energy corrections either using thermally displaced configurations generated by MD, Monte Carlo, or special displacements sampling, yielding renormalized frequencies and linewidths, while preserving the average crystal symmetry and therefore do not explicitly explore local symmetry breaking. A key strength of these methods is the systematic use of crystal symmetries, which accelerates convergence and enables extensions beyond the loop diagram, for example, to capture corrections involving third- and higher-order anharmonicity 42, 32. In contrast, the polymorphous framework focuses on explicitly capturing local symmetry breaking and positional polymorphism. Phonon self energy corrections can be incorporated on top of the polymorphous framework starting from optimized supercells 14.
@@
Figure 3: Impact of local disorder on electron-phonon coupling. (a,b,c) Locally disordered structure and diagrammatic representation of loop diagram contributing to temperature-dependent anharmonicity and the basic electron-phonon coupling process. Lines represent electrons with wavector $\bf k$} and $\bf k+q$} and the zigzag lines the anharmonic phonon. (d,e) Calculated band gap renormalization of monomorphous and polymorphous cubic CsPbI3 and tetragonal/cubic CsPbBr3 by ASDM compared with experiment. Data taken from Refs. 24, 14. (f) Calculated electron and hole effective mass of monomorphous and polymorphous cubic CsPbCl3. Data extracted from Ref. 24. Vertical dashed lines indicate temperatures where phase transitions take place.